# A determinantal formula for orthosymplectic Schur functions

### v1

@@ 3. Determinant formula for orthosymplectic Schur functions
Now applying $C_{j}\rightarrow C_{j}-C_{s}$, for $j\in[1,s-1]$ and then using $h_{p}(y_{j},y_{s+1},y_{s+2},\dots,y_{m})-h_{p}(y_{s},y_{s+1},y_{s+2},\dots,y_{m})=(y_{j}-y_{s})h_{p-1}(y_{j},y_{s},y_{s+1},y_{s+2},\dots,y_{m})$, successively for $s=m-1,m-2,\dots,2$, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}\prod_{q=1}^{j-1}(x_{i}+y_{q})-\bar{x}_{i}\prod_{q=1}^{j-1}(\bar{x}_{i}+y_{q}))\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[2,m]$ and using $h_{n}(y_{j-1},\dots,y_{m})=h_{n}(y_{j},\dots,y_{m})+y_{j-1}h_{n-1}(y_{j-1},\dots,y_{m})$, $\operatorname{spo}_{\lambda}(X/Y)$ is given by
@@
Now again applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[3,m]$, we see that the orthosymplectic Schur function is given by
$\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq 2\end{subarray}}&\left(N_{i,j}^{3}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 3\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq 2\end{subarray}}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-2},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
@@
(3.8) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Observe that the expression in (3.8) is the same as the expression in (3.1). This completes the proof. ∎
@@ 4. A generalization of Brent–Krattenthaler–Warnaar’s identity
In this section, we prove a generalization of the following odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals [BKW16]. Recently, Okada gave its linear-algebraic proof [Oka20].
@@
where $\lambda$ runs over partitions of length at most $n+1$ such that $\lambda_{1}\leq r$, and $(r^{m-n})\cup\lambda$ stands for the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n+1})$.
We write some lemmas before presenting the generalization (see 4.7). Recall the Cauchy–Binet formula. Given an $m\times n$ matrix $X=(x_{i,j})_{{1\leq i\leq m,1\leq j\leq n}}$ and a subset $B\subset[n]=\{1,\dots,n\}$ of column indices, let $X[B]$ be the submatrix of $X$ obtained by picking up colmns indexed by elements of $B$.
@@
where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
@@
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$
@@
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
@@
$\operatorname{spo}_{\lambda}(X/z)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$. Let $A_{\lambda}(X)\coloneqq(a_{i,j})_{1\leq i,j\leq n}$ such that
@@
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)=\frac{\det A_{\lambda}(X)}{x_{1}^{n}\cdots x_{n}^{n}}\times\frac{1}{D}.$
Since $\lambda_{j}+r+2n-j+1>\lambda_{j}+r+2n-j>r-\lambda_{j}+j>r-\lambda_{j}+j-1\geq 0,$ $(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$ is a polynomial in $x_{1},\dots,x_{n}$. Also, note that $r-\lambda_{j}+j-1=0$ if and only if $j=1$ and $\lambda_{1}=r$. Therefore,
$a_{1,j}|_{x_{1}=0}=\begin{cases}-1&\lambda_{1}=r\text{ and }j=1\\ 0&\text{ otherwise.}\end{cases}$
@@
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]|_{x_{1}=0}&=\frac{-(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})}{-(x_{2}\cdots x_{n})^{n}\prod_{i=2}^{n}(x_{i}-\bar{x}_{i})\prod_{2\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}\\ &=(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z).\end{split}$
@@
Let $p(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
@@
Let $q(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
@@
(4.4) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{(r^{m-n})\cup\lambda}&(Y/z)\\ =\sum_{j=1}^{m+n+1}&z^{r+m+n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}(X,Y),\end{split}$
where $\lambda$ runs over partitions of length at most $n$ such that $\lambda_{1}\leq r$.
@@
for $1\leq i\leq n$, $1\leq p\leq M$. Then from a direct but lengthy computation, one can see that the $(i,j)$’th entry of $XY^{\prime}$ is given by
@@
(4.5) $\det XY^{\prime}=(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right).$
@@
for $11\leq j\leq 2n+$ 1. Suppose $I_{n}(\lambda)=(\lambda_{1}+n,\dots,\lambda_{n}+1)$. Then by 1.4 and the Cauchy–Binet formula in 4.2, the left hand side of (4.4) is given by
(4.7) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=&\displaystyle\sum_{\lambda}\frac{z^{r}}{D_{x}D_{y}}\det X[I_{n}(\lambda)]\,{\det Y[I_{n}(\lambda)]}\\ =\frac{z^{r}}{D_{x}D_{y}}\det(XY^{\prime})=&\frac{(-1)^{n}z^{r}(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\times\det W^{(M)},\end{split}$
@@
$\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n}\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W_{0},$
@@
$\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n^{2}+n}}{D_{x}D_{y}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j})}&\det W_{0}\\ =\sum_{j=1}^{2n+1}z^{r+2n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{2n+1-j}\big{)}}&(X,Y),\end{split}$
where the last equality uses (2.11). This completes the proof in the case $m=n$. Now we prove the general case by downward induction on $n$. We assume (4.4). Multiplying both sides of (4.4) by $(x_{1}\cdots x_{n}y_{1}\cdots y_{m})^{r}$ and then substituting $x_{1}=0$, by 4.3 and 4.4, we have
$\begin{split}\sum_{\lambda}z^{r}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)(y_{1}\cdots y_{m})^{r}\operatorname{spo}_{(r^{m-n})\cup\lambda}(Y&/z)\\ =\sum_{j=2}^{m+n+1}z^{r+m+n+1-j}(x_{2}\cdots x_{n}y_{1}\cdots y_{m})^{r}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-2},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
@@
$\begin{split}\sum_{\mu}z^{r}\operatorname{spo}_{\mu}(x_{2},\dots,x_{n}/z)\operatorname{spo}_{(r^{m-n+1})\,\cup\,\mu}(Y&/z)\\ =\sum_{j=1}^{m+n}z^{r+m+n-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$

### v2

@@ 3. Determinant formula for orthosymplectic Schur functions
Now applying the elementary column transformations $C_{j}\rightarrow C_{j}-C_{s}$, $j\in[1,s-1]$ and using $h_{p}(y_{j},y_{s+1},y_{s+2},\dots,y_{m})-h_{p}(y_{s},y_{s+1},y_{s+2},\dots,y_{m})=(y_{j}-y_{s})h_{p-1}(y_{j},y_{s},y_{s+1},y_{s+2},\dots,y_{m}),$ successively for $s=m-1,m-2,\dots,2$, we have
$A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}\prod_{q=1}^{j-1}(x_{i}+y_{q})-\bar{x}_{i}\prod_{q=1}^{j-1}(\bar{x}_{i}+y_{q}))\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Applying the elementary column transformations $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[2,m]$ and using $h_{n}(y_{j-1},\dots,y_{m})=h_{n}(y_{j},\dots,y_{m})+y_{j-1}h_{n-1}(y_{j-1},\dots,y_{m})$, $A$ is given by
@@
Now again applying the elementary column transformations $C_{j}\rightarrow C_{j}+y_{j-2}C_{j-1}$, for $j\in[3,m]$, we see that
$A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq 2\end{subarray}}&\left(N_{i,j}^{3}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 3\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq 2\end{subarray}}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-2},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
@@
(3.10) $A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Observe that the expression in (3.10) is the same as the expression in (3.1), which is equal to $\operatorname{spo}_{\lambda}(X;Y)$. This completes the proof. ∎
@@ 4. A generalization of Brent–Krattenthaler–Warnaar’s identity
In this section, we prove a generalization of the following odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals [BKW16]. Recently, Okada gave a linear-algebraic proof [Oka20].
@@
where $\lambda$ runs over partitions of length at most $n+1$ such that $\lambda_{1}\leq r$, and $(r^{m-n})\cup\lambda$ denotes the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n+1})$.
We recall some lemmas before presenting the generalization (see 4.7). We first recall the Cauchy-Binet formula. Given an $m\times n$ matrix $X=(x_{i,j})_{{1\leq i\leq m,1\leq j\leq n}}$ and a subset $B\subset[n]=\{1,\dots,n\}$ of column indices, let $X[B]$ be the submatrix of $X$ consisting of the columns indexed by elements of $B$.
@@
where $F|_{x_{1}=0}$ is the polynomial obtained by substituting $x_{1}=0$ in $F$.
@@
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)$
@@
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z)&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
where $F|_{x_{1}=0}$ is the polynomial obtained by substituting $x_{1}=0$ in $F$.
@@
$\operatorname{spo}_{\lambda}(X;z)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$. Let $A_{\lambda}(X)\coloneqq(a_{i,j})_{1\leq i,j\leq n},$ where
@@
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)=\frac{\det A_{\lambda}(X)}{x_{1}^{n}\cdots x_{n}^{n}}\times\frac{1}{D}.$
Since
$\lambda_{j}+r+2n-j+1>\lambda_{j}+r+2n-j>r-\lambda_{j}+j>r-\lambda_{j}+j-1\geq 0,$
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)$ is a polynomial in $x_{1},\dots,x_{n}$. Also, note that $r-\lambda_{j}+j-1=0$ if and only if $j=1$ and $\lambda_{1}=r$. Therefore,
$a_{1,j}|_{x_{1}=0}=\begin{cases}-1&\lambda_{1}=r\text{ and }j=1,\\ 0&\text{ otherwise,}\end{cases}$
@@
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)]|_{x_{1}=0}&=\frac{-(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})}{-(x_{2}\cdots x_{n})^{n}\prod_{i=2}^{n}(x_{i}-\bar{x}_{i})\prod_{2\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}\\ &=(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z).\end{split}$
@@
Let $p(x,y,z,a,b)$ be the rational function in $x$, $y$, $z$, $a$ and $b$ given by
@@
Let $q(x,y,z,a,b)$ be the rational function in $x$, $y$, $z$, $a$ and $b$ given by
@@
(4.4) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{(r^{m-n})\cup\lambda}&(Y;z)\\ =\sum_{j=1}^{m+n+1}&z^{r+m+n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}(X,Y),\end{split}$
where $\lambda$ runs over partitions of length at most $n$ such that $\lambda_{1}\leq r$ and $(r^{m-n})\cup\lambda$ denotes the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n})$.
@@
for $1\leq i\leq n$, $1\leq p\leq M$. By using the formula for the geometric sum, it is immediate that the $(i,j)$-th entry of $XY^{t}$ is given by
@@
(4.5) $\det XY^{t}=(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right).$
@@
for $11\leq j\leq 2n+$ 1. Suppose $I_{n}(\lambda)={\{}\lambda_{1}+n,\dots,\lambda_{n}+1{\}}$. Then by 1.4 and the Cauchy–Binet formula in 4.2, the left hand side of (4.4) is given by
(4.7) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)=&\displaystyle\sum_{\lambda}\frac{z^{r}}{D_{x}D_{y}}\det X[I_{n}(\lambda)]\,{\det Y[I_{n}(\lambda)]}\\ =\frac{z^{r}}{D_{x}D_{y}}\det(XY^{t})=&\frac{(-1)^{n}z^{r}(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\times\det W^{(M)},\end{split}$
@@
$\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)=\frac{(-1)^{r+n}\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W_{0},$
@@
$\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)&=\frac{(-1)^{r+n^{2}+n}}{D_{x}D_{y}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j})}\det W_{0}\\ &=\sum_{j=1}^{2n+1}z^{r+2n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{2n+1-j}\big{)}}(X,Y),\end{split}$
where the last equality uses (2.10). This completes the proof in the case $m=n$. Now we prove the general case by downward induction on $n$. Multiplying both sides of (4.4) by $(x_{1}\cdots x_{n}y_{1}\cdots y_{m})^{r}$ and then substituting $x_{1}=0$, by 4.3 and 4.4, we have
$\begin{split}\sum_{\lambda}z^{r}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z)(y_{1}\cdots y_{m})^{r}\operatorname{spo}_{(r^{m-n})\cup\lambda}(Y&;z)\\ =\sum_{j=2}^{m+n+1}z^{r+m+n+1-j}(x_{2}\cdots x_{n}y_{1}\cdots y_{m})^{r}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-2},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
@@
$\begin{split}\sum_{\mu}z^{r}\operatorname{spo}_{\mu}(x_{2},\dots,x_{n};z)&\operatorname{spo}_{(r^{m-n+1})\,\cup\,\mu}(Y;z)\\ =\sum_{j=1}^{m+n}&z^{r+m+n-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n-j}\big{)}}(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
@@
Declarations
The authors have no competing interests to declare that are relevant to the content of this article.